herharley:

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 25 y = 2x − 5

HuskyNation:

x² + y² = 25 y = 2 x - 5


x² + ( 2 x - 5 )² = 25 x² + 4 x² - 20 x + 25 = 25 5 x² - 20 x = 25 - 25 5 x ( x - 4 ) = 0 x 1 = 0 and x 2 = 4 Answer: 4 and 0

HuskyNation:

@herharley

HuskyNation:

hello?

HuskyNation:

@herharley see it? hello

wo1f0mon:

hey no spam @HuskyNation delete all spam plz and you are correct

steve816:

lol come on give him a break, this is hardly considered a spam @wo1f0mon

HuskyNation:

Exactly i was just trying to bring the user back over because i saw her online looking at a different question

Bob:

He's desperate for some mod application it looks like

herharley:

i wasnt onnie thanks for the answer

wo1f0mon:

he had more before @steve816

steve816:

What do you mean he had more? It's not possible to delete posts here currently @wo1f0mon

wo1f0mon:

http://prntscr.com/dy41aw